x^2+3=y
y=5-x
Can someone help me with this i keep getting a negative number !This is substituting Linear quadratic equations.Thanks so much !

Answer by mananth(16949)

x^2+3=y
y=5-x
So x^2+3 = 5-x
x^2+x-2=0
(x+2)(x-1)=0
x=-2 OR x=1